### v2

# Neural Bridge Processes

Abstract
Learning stochastic functions from partially observed context-target pairs requires models that are expressive, uncertainty-aware, and strongly conditioned on inputs. Neural Diffusion Processes (NDPs) improve expressivity with denoising diffusion, but their forward process is input-independent; inputs only enter the reverse denoiser, so the noisy training states themselves do not encode the conditioning inputs. We propose Neural Bridge Processes (NBPs), which replace the unconditional forward kernel with an input-anchored bridge trajectory. When input and output dimensions differ, NBP learns an output-space anchor $a_{\psi}(x)=P_{\psi}(x)$, allowing coordinates or other inputs to guide the generative path without changing the denoising backbone. We show theoretically that process-level anchoring induces pathwise input distinguishability, injects information about $x$ into noisy states, and creates a direct gradient pathway unavailable to NDPs. Experiments on synthetic regression, EEG, CylinderFlow, and image regression show consistent improvements. Additional ablations show that the gains come from the full bridge construction with learned alignment, and that the same input-anchored path principle transfers to Flow Matching Neural Processes. These results suggest that bridge-anchored generative paths provide a general mechanism for strengthening conditional stochastic function modeling.
1 Introduction
Learning stochastic functions from partially observed context-target pairs is a fundamental problem in probabilistic modeling Rasmussen (2003); Garnelo et al. (2018b, a); Dutordoir et al. (2023); Franzese et al. (2023); Bonito et al. (2023); Mathieu et al. (2023); Chang et al. (2024); Dou et al. (2025); Hamad and Rosenbaum , with applications in meta-learning Garnelo et al. (2018b, a), few-shot regression Kim et al. (2019), Bayesian optimization Dutordoir et al. (2023); Krishnamoorthy et al. (2023), and uncertainty-aware prediction Chang et al. (2024); Dou et al. (2025); Hamad and Rosenbaum ; Requeima et al. (2024). These settings require models that generalize across tasks while remaining calibrated under scarce or incomplete observations. Gaussian Processes (GPs) Rasmussen (2003) have traditionally dominated this area due to their analytical tractability and clear uncertainty quantification. However, GPs inherently assume Gaussianity and exhibit cubic computational complexity with respect to data size, severely limiting their applicability in scenarios involving large datasets or inherently non-Gaussian functional distributions Snelson and Ghahramani (2005); Titsias (2009); Xu and Zeng (2024); Xu et al. (2024).
Neural Processes (NPs) Garnelo et al. (2018b, a); Kim et al. (2019); Louizos et al. (2019) offer a scalable alternative by amortizing stochastic function prediction with neural architectures. However, standard NPs often struggle to capture complex multi-modal target distributions, motivating stronger generative mechanisms.
NDPs Dutordoir et al. (2023) address this by modeling the input-output mapping as a learned diffusion process Ho et al. (2020), offering enhanced expressivity and sample diversity. Despite their promise in stochastic function modeling, NDPs rely on an unconditional forward process. The input affects the reverse denoiser, but the noised state $y_{t}$ used during training is sampled from a distribution that does not depend on $x$. This creates a structural gap for conditional function modeling: input information is supplied only through the learned reverse network, rather than through the generative path itself.
In this work, we introduce Neural Bridge Processes (NBP), a diffusion-based generative framework that integrates input supervision throughout the diffusion trajectory. NBPs reformulate the forward diffusion kernel so that intermediate noisy states are dynamically anchored by the input through a bridge coefficient $\gamma_{t}$, together with a reverse correction term that keeps the reverse dynamics consistent with the modified forward process. When input and output dimensions differ, such as in coordinate-to-RGB image regression, we use a learned output-space anchor $a_{\psi}(x)=P_{\psi}(x)$ while leaving the denoising architecture unchanged; details and a visualization are provided in Appendix D.
NBP is related to conditional diffusion and diffusion bridge models, but differs in both the object being modeled and the location of conditioning. It learns distributions over functions and places the input anchor directly in the DDPM-style forward kernel, rather than only conditioning the denoiser or constructing a generic data-to-data bridge Zhou et al. (2023); Yue et al. (2023); Zheng et al. (2024); Li et al. (2023); Peluchetti (2023); He et al. (2024); Shi et al. (2023); Naderiparizi et al. (2025).
We provide a theoretical analysis showing that process-level anchoring gives NBP properties unavailable to standard NDPs: different inputs induce different forward marginals, the noisy state carries nonzero information about the input under a Gaussian input model, and the denoising objective receives an additional direct gradient pathway through $y_{t}$. We validate NBPs on synthetic data, electroencephalogram (EEG) regression Zhang et al. (1995), CylinderFlow, and image regression. Appendix H.2 isolates the source of the gains through component ablations, and Appendix H.3 shows that the same bridge idea transfers to Flow Matching Neural Processes (FlowNP) Hamad and Rosenbaum .
In summary, the core contributions of this paper are:
  • We introduce Neural Bridge Processes (NBPs), a class of stochastic function models that use input-anchored diffusion trajectories to make the noisy training states explicitly depend on the conditioning inputs.
  • We formulate a DDPM-style bridge for function modeling, including a learned output-space anchor $a_{\psi}(x)=P_{\psi}(x)$ for dimension-mismatched inputs and outputs.
  • We provide theoretical analysis showing that process-level anchoring induces pathwise input distinguishability, input information in noisy states, and a direct gradient pathway for supervision.
  • We validate NBPs on synthetic regression, EEG, CylinderFlow, and image regression, and provide ablations showing that the gains come from the bridge construction rather than merely architecture size or denoiser capacity. We further show that the same bridge idea can improve a Flow Matching Neural Process, suggesting broader relevance for generative NP models.
2 Related Work
Neural Processes (NPs) Garnelo et al. (2018b, a); Kim et al. (2019); Louizos et al. (2019) learn distributions over functions from context-target observations, combining amortized neural prediction with uncertainty modeling. A broad line of work has improved NP expressivity through attention, convolutional structure, autoregressive modeling, graph structure, and other process-aware architectures Gordon et al. (2019); Foong et al. (2020); Singh et al. (2019); Nguyen and Grover (2022); Bruinsma et al. (2023); Hu et al. (2023). More recent methods incorporate stronger generative mechanisms into function modeling, including diffusion-based neural processes Dutordoir et al. (2023); Mathieu et al. (2023), score-based neural processes Dou et al. (2025), flow-based models Hamad and Rosenbaum , and related neural ODE formulations Chen et al. (2018); Norcliffe et al. (2021).
Our work is closest to Neural Diffusion Processes (NDPs) Dutordoir et al. (2023), which are already a conditional diffusion baseline for Neural Process-style function learning: they learn a DDPM-style reverse denoiser conditioned on the input and context observations. The key difference is that NDPs retain an unconditional forward diffusion process, so the noised training states are generated independently of the input. NBP keeps the same conditional denoising backbone but modifies the forward transition itself, so the input acts as a dynamic anchor throughout the diffusion trajectory. This process-level modification distinguishes NBP from standard conditional diffusion models Choi et al. (2021); Zhang et al. (2023); Zhu et al. (2023), where conditioning typically enters the reverse model while the forward noising process remains condition-independent.
NBP is also related to diffusion bridge models Zhou et al. (2023); Yue et al. (2023); Zheng et al. (2024); Li et al. (2023); Peluchetti (2023); He et al. (2024); Shi et al. (2023); Naderiparizi et al. (2025). Diffusion bridges have become popular in generative modeling, but existing DDBM-style methods are usually designed for data-space transport between paired endpoint distributions, such as image-to-image translation. This differs from the Neural Process setting, where the goal is to model stochastic functions from context-target observations and predict outputs at arbitrary target inputs. To the best of our knowledge, NBP is the first to introduce bridge-anchored generative paths into Neural Process-style function learning. The endpoint anchor is the function input, or a learned projection of it, and the bridge is used to strengthen conditional function prediction rather than generic domain translation. Appendix H.3 further compares this bridge principle with recent flow-based Neural Processes.
3 Preliminaries
Neural Processes (NPs) approximate a stochastic process $F:X\to Y$ through finite-dimensional marginals. Given context observations $(x_{\mathcal{C}},y_{\mathcal{C}})$ and target inputs $x_{\mathcal{T}}$, a latent-variable NP predicts target outputs by
$p(y_{\mathcal{T}},z|x_{\mathcal{T}},x_{\mathcal{C}},y_{\mathcal{C}})=p(z|x_{\mathcal{C}},y_{\mathcal{C}})\prod_{i=1}^{|\mathcal{T}|}p(y_{\mathcal{T},i}|x_{\mathcal{T},i},z)$ (1)
where $z$ captures uncertainty about the global structure of the underlying function. NDPs replace simple latent-variable decoders with a learned diffusion process over outputs, which improves expressivity but still uses an unconditional forward noising process.
3.1 Neural Diffusion Processes
Given clean outputs $y_{0}$, NDPs Dutordoir et al. (2023) define the standard DDPM forward process
$q(y_{1:T}\mid y_{0})=$ $\prod_{t=1}^{T}q(y_{t}\mid y_{t-1}),$ (2)
$q(y_{t}\mid y_{t-1})=$ $\mathcal{N}(y_{t};\sqrt{1-\beta_{t}}\,y_{t-1},\beta_{t}I).$
The terminal variable is therefore approximately standard Gaussian noise and does not encode the conditioning input. The reverse process is conditioned on $x$ through a denoising network:
$p_{\theta}(y_{0:T}\mid x)=p(y_{T})\prod_{t=1}^{T}p_{\theta}(y_{t-1}\mid y_{t},x),$ (3)
where $p_{\theta}(y_{t-1}\mid y_{t},x)$ is parameterized by a noise predictor $\epsilon_{\theta}(y_{t},t,x)$. Training uses the usual denoising objective
$\mathcal{L}_{\theta}=$ $\mathbb{E}_{t,x,y_{0},\epsilon}\left[\left\|\epsilon-\epsilon_{\theta}(y_{t},t,x)\right\|_{2}^{2}\right],$ (4)
$\text{with}\quad y_{t}=$ $\sqrt{\bar{\alpha}_{t}}y_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon.$
This compact review highlights the key design choice that motivates NBP: in NDPs, $x$ affects the reverse denoiser but not the forward trajectory.
4 Method: Neural Bridge Processes (NBP)
4.1 Problem Setup
We consider the standard meta-learning setting where a model observes a set of context points $x_{\mathcal{C}}=\{(x_{i},y_{i})\}_{i=1}^{N_{c}}$ and aims to predict the corresponding outputs $y_{\mathcal{T}}=\{y_{j}\}_{j=1}^{N_{t}}$ for a set of target inputs $x_{\mathcal{T}}=\{x_{j}\}_{j=1}^{N_{t}}$. Here, each task is assumed to be sampled from a distribution over functions, and the goal is to model the conditional distribution $p(y_{\mathcal{T}}|x_{\mathcal{C}},x_{\mathcal{T}},y_{\mathcal{C}})$. Throughout the method section, we use $x$ and $y$ to denote the full set of inputs and outputs for one task-level function, while $(x_{\mathcal{C}},y_{\mathcal{C}})$ and $(x_{\mathcal{T}},y_{\mathcal{T}})$ denote the context and target subsets used for conditional prediction. We use calligraphic symbols $\mathcal{C}$ and $\mathcal{T}$ for context and target index sets to avoid confusion with the complex numbers. The diffusion process is defined on the full task-level output set $y_{0:T}$ conditioned on the corresponding inputs $x$; the context–target partition is then instantiated during conditional sampling.
4.2 Motivation
Traditional Neural Diffusion Processes (NDPs) condition on inputs through the denoising network, for example by concatenation or attention. The forward process itself remains an unconditional Gaussian noising process, so the noisy state $y_{t}$ used for training is generated without direct dependence on $x$. By “indirect”, we mean that $x$ influences the model only through the learned denoiser $\epsilon_{\theta}(y_{t},x,t)$, while the distribution that produces $y_{t}$ is independent of $x$. Here, “temporal” refers to the ordered diffusion trajectory indexed by timestep $t$, not to an assumption that NP tasks themselves must have temporal dynamics. Once a diffusion-based NP model is adopted, there is an additional design choice: whether the input-output relation is imposed only through reverse-time conditioning, or also through the forward path used to generate training states. Standard NDPs choose the former.
NBP addresses this mismatch by making the input part of the forward transition. Instead of treating $x$ only as side information for the denoiser, we use $x$ as a dynamic anchor that changes the mean of every noising step. This coupling between $x$ and $y_{t}$ is therefore present throughout the trajectory rather than only in the denoiser. The resulting process remains DDPM-style and easy to train, but its trajectory is explicitly input-aligned in expectation.
To make this motivation concrete, we use Jacobian sensitivity as a diagnostic of effective input coupling:
$S(t)=\mathbb{E}\left[\left\|\frac{\partial\epsilon_{\theta}(y_{t},x,t)}{\partial x}\right\|_{F}\right].$ (5)
We compute this quantity on CelebA $32\times 32$ with context ratio 0.3 by averaging over 100 test images and multiple diffusion timesteps; implementation details are provided in Appendix H.4. Figure 1 shows that NDP has lower sensitivity to $x$ at early timesteps, whereas NBP has a larger measured sensitivity in the same region because $x$ also enters the forward state through the bridge term. We use this figure as supporting diagnostic evidence, not as a standalone proof of performance: the main quantitative evidence is provided by the likelihood, error, calibration, and ablation results.
Figure 1: Jacobian sensitivity across diffusion timesteps on CelebA $32\times 32$ with context ratio 0.3, averaged over 100 test images. NBP shows larger early-timestep sensitivity than NDP.
4.3 Bridge Construction
We first present the construction when the input and output share the same dimensionality. Given observed pairs $(x,y)$ with $x,y\in\mathbb{R}^{N\times D}$ and clean output $y_{0}=y$, NBP replaces the unconditional DDPM forward transition with
$q(y_{t}|y_{t-1},x)=\mathcal{N}\left(y_{t};\underbrace{\sqrt{1-\beta_{t}}y_{t-1}+\gamma_{t}x}_{\text{Bridge-anchored mean}},\beta_{t}I\right).$ (6)
The scalar $\gamma_{t}$ controls the strength of the input anchor. In this sense, the anchor is dynamic: the bridge term $\gamma_{t}x$ changes with the diffusion timestep and biases each intermediate state toward the input-conditioned endpoint with time-varying strength. We use the SNR-aware schedule
$\gamma_{t}=\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}},\quad\mathrm{SNR}_{t}=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}.$ (7)
When the input and output dimensions differ, we replace the raw input anchor $x$ by a learned output-space anchor $a_{\psi}(x)=P_{\psi}(x)\in\mathbb{R}^{N\times D_{y}}$ and apply the same bridge construction to $a_{\psi}(x)$. For example, in image regression, $x$ is a two-dimensional spatial coordinate while $y$ is an RGB value; the learned anchor network maps coordinates to the output channel space before they enter the bridge. Appendix D and Figure 3 provide the dimensional-alignment details. For readability, the equations below use $x$ to denote the aligned anchor, with the general case obtained by substituting $x$ with $a_{\psi}(x)$. This schedule keeps the early trajectory close to standard diffusion, where $\mathrm{SNR}_{t}$ is high and $\gamma_{t}$ is small, while increasing input attraction near the terminal region. Since $\gamma_{T}=\mathrm{SNR}_{T}/\mathrm{SNR}_{T}=1$, the last single-step transition has full input anchoring. This does not imply that the cumulative coefficient $\bar{\gamma}_{T}$ equals one; its scale is determined by the entire noise schedule. The following proposition gives the closed-form marginal used for training.
Proposition 4.1 (Input-anchored forward marginal).
For the transition in Eq. (6), the marginal distribution at any timestep $t$ is
$y_{t}\mid y_{0},x\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\,y_{0}+\bar{\gamma}_{t}x,\;(1-\bar{\alpha}_{t})I\right),$ (8)
where
$\bar{\gamma}_{t}=\sum_{s=1}^{t}\gamma_{s}\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}}.$ (9)
The proof is provided in Appendix B.1. Proposition 4.1 shows that the terminal distribution is no longer input-independent. When $\bar{\alpha}_{T}\approx 0$, the terminal mean satisfies
$\mathbb{E}[y_{T}|y_{0},x]=\sqrt{\bar{\alpha}_{T}}y_{0}+\bar{\gamma}_{T}x\approx\bar{\gamma}_{T}x,$ (10)
while the terminal covariance is exactly $(1-\bar{\alpha}_{T})I$, which approaches $I$ only under a sufficiently long or strong noise schedule. Thus, the terminal state remains stochastic and schedule-dependent, but its mean becomes aligned with the input anchor in expectation.
Reverse Process with Bridge Correction
The reverse model keeps the DDPM denoising parameterization but adds the correction term induced by the input-dependent forward kernel:
$p_{\theta}(y_{t-1}|y_{t},x)=\mathcal{N}\left(y_{t-1};\mu_{\theta}(y_{t},x,t),\tilde{\beta}_{t}I\right)$ (11)
where $\alpha_{t}=1-\beta_{t}$, and
$\mu_{\theta}=\underbrace{\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(y_{t},x,t)\right)}_{\text{Denoising term}}+\underbrace{C_{t}(x)}_{\text{Bridge correction}}$ (12)
Proposition 4.2 (Bridge posterior correction).
The posterior mean of $q(y_{t-1}\mid y_{t},y_{0},x)$ can be written in DDPM noise-prediction form plus the bridge correction
$C_{t}(x)=-\frac{\gamma_{t}}{\sqrt{1-\beta_{t}}}x=-\frac{\gamma_{t}}{\sqrt{\alpha_{t}}}x.$ (13)
The derivation is given in Appendix B.2. This correction is not an additional heuristic; it follows from the Gaussian posterior induced by the bridge transition in Eq. (6). Intuitively, the forward step injects $\gamma_{t}x$ into the mean, and the reverse mean must compensate for this deterministic input shift while denoising the stochastic component. This keeps the learned reverse process consistent with the bridge-anchored forward dynamics.
Training Objective
Proposition 4.3 (Denoising objective).
Optimizing the variational bound for the bridge process reduces, up to the standard DDPM weighting, to the noise-prediction objective
$\mathcal{L}_{\theta}=$ $\mathbb{E}_{t,x,y_{0},\epsilon}\left[\|\epsilon_{\theta}(y_{t},x,t)-\epsilon\|_{2}^{2}\right]$ (14)
$\text{with}\quad y_{t}=$ $\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon.$
The proof is provided in Appendix B.3. Importantly, NBP does not require a new denoising architecture: the same noise predictor is trained with bridge-anchored samples from Proposition 4.1.
Conditional Sampling Procedure.
At test time, NBPs sample from $p(y_{0}\mid x,\mathcal{D})$, where $\mathcal{D}=(x_{\mathcal{C}},y_{\mathcal{C},0})$ denotes the observed context set. We initialize the terminal state from the input-aligned bridge endpoint,
$y_{T}=\bar{\gamma}_{T}x+n,\quad n\sim\mathcal{N}(0,(1-\bar{\alpha}_{T})I).$ (15)
For each reverse timestep, the observed context outputs are re-noised using the closed-form bridge marginal
$y_{\mathcal{C},t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}y_{\mathcal{C},0}+\bar{\gamma}_{t}a_{\psi}(x_{\mathcal{C}}),(1-\bar{\alpha}_{t})I\right),$ (16)
then combined with the current target state as $y_{t}=\{y_{\mathcal{T},t},y_{\mathcal{C},t}\}$ and $x=\{x_{\mathcal{T}},x_{\mathcal{C}}\}$. The reverse transition is then sampled from
$y_{t-1}\sim\mathcal{N}\left(\mu_{\theta}(y_{t},x,t),\tilde{\beta}_{t}I\right),$ (17)
where $\mu_{\theta}$ includes the bridge correction in Eq. (13). Following RePaint Lugmayr et al. (2022), we may repeat the context re-noising and reverse denoising step at each timestep before moving to t1t-1. This repeatedly reinjects the observed context and improves conditional coherence along the generated trajectory.
4.4 Theoretical Analysis: Why Process-Level Anchoring Helps
The preceding propositions establish that NBP is a valid DDPM-style bridge construction. We next make explicit what is gained by placing the input in the forward process rather than only in the denoising network. The goal is not to claim a universal generalization guarantee, but to identify structural effects that follow directly from the bridge kernel.
Theorem 4.4 (Pathwise input identifiability).
Fix $y_{0}$ and consider two inputs $x,x^{\prime}$. Let $q_{t}^{\mathrm{NBP}}(\cdot|y_{0},x)$ denote the NBP forward marginal in Eq. (8). Then
$W_{2}\!\left(q_{t}^{\mathrm{NBP}}(\cdot|y_{0},x),q_{t}^{\mathrm{NBP}}(\cdot|y_{0},x^{\prime})\right)=|\bar{\gamma}_{t}|\,\|x-x^{\prime}\|_{2}.$ (18)
By contrast, the NDP forward marginal is independent of $x$, and therefore the corresponding Wasserstein distance is zero for all $x,x^{\prime}$.
Theorem 4.4 formalizes the sense in which NBP creates pathwise input coupling: different inputs induce different noisy states along the forward trajectory. A larger conditioning network can make the reverse model more expressive, but it does not change the fact that the NDP training states are generated from an input-independent forward kernel.
Theorem 4.5 (Forward-state information about the input).
Assume $x\sim\mathcal{N}(0,\Sigma_{x})$ and condition on $y_{0}$. Under the NBP forward marginal,
$I(x;y_{t}\mid y_{0})=\frac{1}{2}\log\det\!\left(I+\frac{\bar{\gamma}_{t}^{2}}{1-\bar{\alpha}_{t}}\Sigma_{x}\right).$ (19)
For the NDP forward process, $I(x;y_{t}\mid y_{0})=0$ because $y_{t}$ is conditionally independent of $x$ given $y_{0}$.
This gives an information-theoretic interpretation of the bridge term. NBP makes the noisy training state itself informative about the input; NDP leaves all input information to the denoiser-side conditioning pathway. The amount of information depends on the schedule through the signal-to-noise ratio $\bar{\gamma}_{t}^{2}/(1-\bar{\alpha}_{t})$, which also motivates schedule selection.
Proposition 4.6 (Direct gradient pathway).
For a single-timestep denoising loss
$\ell_{t}(\theta,x)$ $=\|\epsilon_{\theta}(y_{t},x,t)-\epsilon\|_{2}^{2},$ (20)
$y_{t}$ $=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$
the total derivative with respect to $x$ decomposes as
$\frac{d\ell_{t}}{dx}=\underbrace{\frac{\partial\ell_{t}}{\partial x}}_{\text{denoiser conditioning path}}+\underbrace{\bar{\gamma}_{t}\frac{\partial\ell_{t}}{\partial y_{t}}}_{\text{forward bridge path}}.$ (21)
For NDP, $\bar{\gamma}_{t}=0$ in the forward process, so only the denoiser conditioning path remains.
Proposition 4.6 explains why the bridge construction can improve training even when the denoising architecture is unchanged. NBP supplies an additional route by which input supervision affects the denoising loss through the sampled state $y_{t}$ itself. This complements the empirical Jacobian sensitivity in Figure 1, where NBP shows stronger input sensitivity especially in early diffusion steps.
Table 1: Mean test log-likelihood ($\uparrow$) $\pm$ 1 standard error estimated over 128 test samples.
Squared Exponential Matérn-$\frac{5}{2}$
Model $D=1$ $D=2$ $D=3$ $D=1$ $D=2$ $D=3$
ANP $4.79{\pm}0.05$ $23.80{\pm}0.05$ $23.80{\pm}0.04$ $0.70{\pm}0.04$ $17.22{\pm}0.02$ $21.24{\pm}0.02$
ConvCNP $6.40{\pm}0.07$ $24.00{\pm}0.03$ $23.80{\pm}0.02$ $0.87{\pm}0.06$ $17.50{\pm}0.03$ $21.24{\pm}0.02$
GNP $4.00{\pm}0.02$ $19.60{\pm}0.02$ $23.80{\pm}0.02$ $\mathbf{0.14}{\pm}0.02$ $15.70{\pm}0.02$ $21.20{\pm}0.02$
NDP $4.21{\pm}0.04$ $13.39{\pm}0.05$ $20.48{\pm}0.05$ $0.13{\pm}0.02$ $14.74{\pm}0.03$ $20.66{\pm}0.05$
SNP $4.27{\pm}0.02$ $13.19{\pm}0.03$ $20.24{\pm}0.04$ $0.01{\pm}0.02$ $14.67{\pm}0.03$ $20.59{\pm}0.05$
GEOMNDP $4.22{\pm}0.05$ $13.36{\pm}0.05$ $20.45{\pm}0.05$ $0.13{\pm}0.03$ $14.73{\pm}0.03$ $20.63{\pm}0.06$
NBP (ours) $\mathbf{4.72}{\pm}0.03$ $\mathbf{-12.10}{\pm}0.05$ $\mathbf{-18.10}{\pm}0.04$ $0.047{\pm}0.019$ $\mathbf{-13.30}{\pm}0.03$ $\mathbf{-18.46}{\pm}0.03$
Table 2: Predictive NLL ($\downarrow$) and MSE ($\downarrow$) on EEG
Method Interpolation Reconstruction Forecasting
NLL MSE($\times 10^{-2}$) NLL MSE($\times 10^{-2}$) NLL MSE($\times 10^{-2}$)
NP $1.66\pm 0.08$ $0.52\pm 0.03$ $1.78\pm 0.09$ $0.44\pm 0.02$ $1.61\pm 0.07$ $0.39\pm 0.02$
ANP $0.47\pm 0.05$ $0.25\pm 0.02$ $0.70\pm 0.06$ $0.48\pm 0.03$ $0.90\pm 0.06$ $0.60\pm 0.04$
ConvNP $0.44\pm 0.04$ $0.40\pm 0.03$ $2.43\pm 0.12$ $0.40\pm 0.02$ $2.34\pm 0.11$ $0.55\pm 0.03$
NDP $2.46\pm 0.11$ $0.18\pm 0.01$ $2.59\pm 0.13$ $0.23\pm 0.02$ $2.69\pm 0.12$ $0.38\pm 0.02$
SNP $3.19\pm 0.10$ $0.16\pm 0.01$ $\mathbf{-3.30\pm 0.11}$ $0.18\pm 0.01$ $3.02\pm 0.09$ $0.31\pm 0.02$
GEOMNDP $2.48\pm 0.12$ $0.18\pm 0.01$ $2.65\pm 0.13$ $0.20\pm 0.01$ $2.84\pm 0.11$ $0.34\pm 0.02$
NBP (Ours) $\mathbf{-3.55\pm 0.10}$ $\mathbf{0.15\pm 0.01}$ $\mathbf{-3.45\pm 0.11}$ $\mathbf{0.15\pm 0.01}$ $\mathbf{-3.74\pm 0.11}$ $\mathbf{0.27\pm 0.02}$
5 Experiments
5.1 Baseline Implementation and Evaluation Metrics
For a comprehensive comparison, we implement Neural Processes (NPs) Garnelo et al. (2018b), Attentive Neural Processes (ANPs) Kim et al. (2019), and Convolutional Neural Processes (ConvNPs) Gordon et al. (2019) using the official NP-Family repository Dubois et al. (2020), with all hyperparameters set to the recommended defaults. We further include Gaussian Neural Processes (GNPs) Bruinsma et al. (2021) in our synthetic experiments. For Neural Diffusion Processes (NDPs) Dutordoir et al. (2023), Geometric Neural Diffusion Processes (GEOMNDPs) Mathieu et al. (2023), and Score-Based Neural Processes (SNPs) Dou et al. (2025), we directly adopt their official implementations. To ensure a controlled comparison with NDP, our Neural Bridge Process (NBP) uses the same Bi-Dimensional Attention Block denoising backbone, optimizer, diffusion schedule, number of denoising steps, and training budget as the NDP baseline. The bridge coefficients and reverse correction introduce no trainable parameters. When $D_{x}\neq D_{y}$, NBP additionally uses the learned anchor map $P_{\psi}$, which is part of the proposed bridge construction rather than a replacement for the denoiser; we therefore report fixed-anchor and learned-anchor ablations in Appendix H.2 to separate this effect from the bridge itself. Detailed implementation settings are provided in Appendix G. Additional ablation studies, including bridge coefficient schedules, component ablations on CelebA NLL, a bridge-transfer experiment on FlowNP, and an analysis quantifying path supervision, are presented in Appendices H, H.2, and H.3. All models are retrained on the respective experimental datasets and evaluated with the same set of random seeds; reported values are averages over multiple seeds unless otherwise specified. Experiments are conducted on NVIDIA RTX A6000 GPUs with 48GB memory; each reported run uses a single GPU unless otherwise stated.
5.2 Regression on Synthetic Data
We evaluate our method on synthetic 1D–3D regression tasks, using functions sampled from Gaussian Processes (GPs) with either a Squared Exponential or Matérn-5/2 kernel. For each dimension $D$, the kernel lengthscale is set to $\ell=\sqrt{D}/4$, and Gaussian noise $\mathcal{N}(0,0.05^{2})$ is added to the outputs. During training, we generate $2^{10}$ examples per epoch, and train for 400 epochs using batch size 32. Each model is trained with its own architecture and optimization settings (details below). At test time, the context set contains a random number of points between $1$ and $10\times D$, while the target set always includes 50 points.
The log-likelihood is estimated by fitting a multivariate Gaussian to 128 samples drawn from the conditional distribution of the model. For our proposed method, we use a 4-layer transformer-style architecture with 8 attention heads and 64-dimensional hidden layers. Diffusion noise is scheduled over 500 timesteps with a cosine schedule ($\beta\in[3\mathrm{e}{-4},0.5]$). The optimizer uses a peak learning rate of $10^{-3}$ with warm-up (20 epochs) and cosine decay (200 epochs). All experiments use the same evaluation batch size and sampling procedure for consistency. When input and output dimensions differ, the bridge uses the learned output-space anchor $a_{\psi}(x)=P_{\psi}(x)$ described in Appendix D; the raw input $x$ is still provided to the denoising network $\epsilon_{\theta}$ as conditioning information.
Table 1 shows that NBP consistently outperforms prior Neural Process variants across all input dimensions. Notably, NBP maintains stable and accurate predictions in higher dimensions ($D=2,3$), where performance of other models tends to degrade sharply.
5.3 Real World Experiments
5.3.1 EEG Signal Regression Tasks
We evaluate the proposed Neural Bridge Processes (NBP) on a real-world electroencephalogram (EEG) dataset regression task Zhang et al. (1995). The dataset comprises 7632 multivariate time series, each consisting of 256 evenly sampled time steps recorded across seven electrode channels. These EEG signals show strong temporal dynamics and cross-channel correlations, making them ideal for evaluating multi-output meta-learning models like NBPs.
To assess NBPs on correlated multi-output prediction and missing data, we randomly mask windows in 3 of 7 channels and predict the missing values. Inputs are represented by the concatenated temporal and channel indices $\mathbf{x}_{e}=(i_{t},i_{c})$, and outputs are the corresponding voltage measurements $\mathbf{y}_{e}$. We evaluate three conditional prediction regimes in a unified setting. In interpolation, the model recovers locally missing values within the observed temporal span. In reconstruction, the model infers partially obscured temporal segments of a target channel using the remaining observed channels as context. In forecasting, the model extrapolates future signal trajectories from current observations. These regimes test complementary aspects of conditional function modeling: local completion, cross-channel reconstruction, and temporal extrapolation. We report Mean Squared Error (MSE) and Negative Log-Likelihood (NLL). As shown in Table 2, NBP consistently outperforms the baselines across the three settings, indicating that the bridge-anchored diffusion process can capture both temporal structure and cross-channel dependencies in EEG signals. Appendix Table 3 further reports calibration metrics on the interpolation setting, where NBP achieves the lowest ECE and sharpest predictive uncertainty while maintaining coverage close to the nominal 90% level.
CylinderFlow.
To further evaluate the empirical scope of NBP beyond synthetic functions and EEG signals, we also report results on the CylinderFlow benchmark, which requires structured spatial prediction under partial observations. Table 5 shows that NBP achieves the lowest MSE among all compared methods, improving over both classical NP variants and the diffusion-based NDP baseline. This suggests that the benefit of bridge-anchored diffusion generalizes to real-world spatial prediction tasks.
Table 4: MSE ($\downarrow$) on the CylinderFlow benchmark.
Method MSE
NP 0.0108
ANP 0.0157
ConvNP 0.0341
NDP 0.0053
NBP (Ours) 0.0031
Table 5: CelebA image regression NLL ($\downarrow$) under half-image and random-context observations.
Method Half Random
NP -0.84 -1.07
ANP -1.39 -1.43
ConvNP 0.11 -1.67
NDP -3.92 -3.98
SNP -4.76 -4.77
NBP (Ours) -5.02 -5.24
5.3.2 Image Regression Task
Figure 2: Comparison of reconstruction errors (MSE) between the NDP Base and our proposed NBP on the CelebA dataset at resolutions of 32×32 and 64×64. The horizontal axis represents the context ratio (i.e., the proportion of retained pixels), while the vertical axis shows the reconstruction error in units of $10^{-2}$. Solid lines indicate the mean MSE across test samples, and the shaded regions represent the standard deviation (Std), reflecting model uncertainty.
We evaluate image regression on CelebA at resolutions of $32\times 32$ and $64\times 64$, where models predict pixel values from spatial coordinates normalized to [2,2][-2,2]. Hyperparameters, denoising architecture, learning rate, seed set, and baseline configurations follow the NDP setup for a controlled comparison. Our baselines focus on NP-family and diffusion/score-based Neural Process models, which share the same conditional function-learning setting.
Figure 2 summarizes MSE across context ratios, with numerical values in Appendix Tables 7 and 8 and qualitative examples in Appendix Figure 6. NBP consistently improves over the NDP baseline and has smaller standard deviations across context ratios. All MSE values average nine conditional samples per input, with pixel values normalized to $[0,1]$. Table 5 further shows NLL improvements over broader NP-family baselines, while Appendix H.2 decomposes the gain into endpoint initialization, bridge construction, anchor learning, and denoiser-capacity effects.
For example, at context ratio $0.2$, NBP reduces MSE from $0.88$ to $0.76$ on CelebA $32\times 32$ and from $1.05$ to $0.86$ on CelebA $64\times 64$, supporting the usefulness of bridge-anchored trajectories for sparse-coordinate image regression.
5.4 Computational Efficiency
Under identical training settings, NBP and the NDP baseline exhibit comparable training time per epoch and overall runtime. NBP uses the same denoising architecture and number of sampling steps as NDP; the additional operations are limited to schedule-dependent bridge coefficients and a simple correction term. Thus, the practical overhead is negligible in our implementation, while the improvement comes from the process-level input–output coupling rather than additional network capacity.
6 Conclusion
In this work, we introduced Neural Bridge Processes (NBPs), a diffusion-based framework for stochastic function modeling that explicitly incorporates input supervision throughout the diffusion trajectory. By reformulating the forward kernel with a bridge coefficient and, when needed, a learned output-space anchor, NBPs address the weak input coupling and endpoint mismatch of traditional Neural Diffusion Processes (NDPs) while preserving the same denoising backbone. Experiments on synthetic data, EEG time series, CylinderFlow, and image regression show improved predictive accuracy and uncertainty calibration. The component ablations and FlowNP transfer study further suggest that the benefit comes from input-anchored generative paths rather than only additional capacity or a DDPM-specific implementation. These results highlight a broader possibility: bridge-anchored paths may be a useful design principle for generative Neural Processes, including diffusion-, score-, and flow-based variants.
Appendix A Additional Related Discussion
A.1 Neural Processes and Their Extensions
Neural Processes (NPs) [14] combine the flexibility of neural networks with the uncertainty modeling capabilities of stochastic processes, aiming to learn distributions over functions. Conditional Neural Processes (CNPs) [13] extend this framework by conditioning on observed context points to predict target outputs. Attentive Neural Processes (ANPs) [23] further enhance NPs by incorporating attention mechanisms, enabling the model to focus on relevant context points for each target prediction. Despite these advancements, challenges remain in capturing complex, multimodal distributions and ensuring consistency in posterior predictions. These approaches have been successfully extended to various domains, including sequential modeling [41, 32, 3], convolutional architectures [16, 11], graph-based models [20], and probabilistic predictive models for large language models (LLMs) [38].
A.2 Diffusion, Conditional Diffusion, and Bridge Models
Denoising Diffusion Probabilistic Models (DDPMs) [19] are powerful generative models that approximate complex data distributions by reversing a progressive noising process. Conditional Diffusion Models (CDMs) [7, 50, 54] extend this framework by incorporating auxiliary information, enabling conditional generation. In most CDMs, however, the forward noising process remains unconditional and the condition is injected into the reverse denoising network. This makes CDMs broadly applicable, but it does not explicitly define a condition-dependent forward trajectory or a condition-aligned terminal distribution. NBP differs from this design choice: the input $x$ appears directly in the forward transition kernel, so conditioning affects the entire diffusion path rather than only the learned denoiser.
More recently, Denoising Diffusion Bridge Models (DDBMs) [53, 49, 52, 25, 36, 18, 40, 31] have been proposed as a natural alternative. DDBMs introduce diffusion bridges—stochastic processes that interpolate between two paired data distributions given as endpoints—making them well-suited for tasks such as image-to-image translation. NBP shares the general bridge intuition of endpoint-aware generative paths, but its target is different. Rather than learning a generic data-space transport between two domains, NBP is designed for stochastic function modeling from context-target observations. The endpoint anchor is the function input, or its learned projection when dimensions differ, and the bridge is used to strengthen conditional function prediction. To the best of our knowledge, this is the first use of bridge-anchored generative paths as a Neural Process mechanism.
Thus, NBP should be viewed neither as a standard CDM nor as a direct application of existing DDBMs. It is a DDPM-style bridge formulation specialized to Neural Process-style function learning: it preserves the architecture and training protocol of the conditional diffusion baseline NDP while modifying the process-level forward kernel and corresponding reverse correction. Since DDBMs address a different problem setting, a direct experimental comparison with the original DDBM is not the main focus; instead, we compare against diffusion-based function modeling baselines such as NDP, Geometric NDP, and SNP, and use component ablations to rule out simpler explanations such as endpoint initialization, coordinate-anchor capacity, or a stronger denoiser.
A.3 Generative Models for Function Modeling
Recent work has explored the use of diffusion models for function modeling. Neural Diffusion Processes (NDPs) [10] model distributions over functions by applying diffusion processes in latent space, allowing the representation of complex, non-Gaussian function distributions. Geometric Neural Diffusion Processes [30] further extend this approach by incorporating geometric priors for infinite-dimensional modeling in non-Euclidean spaces. In parallel, other generative modeling techniques such as Neural ODEs [6, 34], flow matching [26, 17], and score-based SDE methods [44, 8] are also being integrated into the Neural Processes (NP) framework to enhance function modeling capabilities. In our experiments, we compared against open-source methods, including score-based neural processes (SNP) [8] and Geometric Neural Diffusion Processes [30], both of which demonstrated the empirical advantages of our approach.
A.4 Flow Matching Neural Processes
Flow Matching Neural Processes (FlowNP) [17] provide a recent flow-based alternative to diffusion-style Neural Processes. Rather than learning a denoising score or noise predictor over a discrete diffusion chain, FlowNP defines a continuous conditional probability path between a simple base sample $y_{0}$ and the observed target value $y_{1}$, commonly using a linear interpolation of the form $y_{t}=(1-t)y_{0}+ty_{1}$. A transformer-based Neural Process backbone then predicts the corresponding velocity field conditioned on the context set. At test time, conditional samples are obtained by integrating the learned vector field, and likelihoods can be estimated through the ODE change-of-variables formula.
This perspective is useful for two reasons. First, FlowNP is a strong and recent baseline for stochastic function modeling, so comparisons against it help clarify whether improvements come from the bridge construction rather than from using a stronger backbone. Second, its path-based formulation is closely related to our motivation: both approaches view conditional function prediction as learning a generative trajectory over function values. The key distinction is that FlowNP uses an input-independent interpolation path and injects context through the velocity network, whereas NBP modifies the generative path itself by anchoring intermediate states to the input. In additional controlled experiments, we therefore also consider bridge-augmented flow paths using the same FlowNP architecture and likelihood evaluation protocol.
Appendix B Proofs for Neural Bridge Processes
B.1 Proof of Proposition 4.1
We derive the closed-form marginal distribution in Proposition 4.1. The forward process depends on the input anchor $x$.
From $y_{t-1}$ to $y_{t}$:
$y_{t}=\sqrt{1-\beta_{t}}y_{t-1}+\gamma_{t}x+\sqrt{\beta_{t}}\epsilon_{t}$ (22)
By recursively expanding:
$y_{t}$ $=\sqrt{1-\beta_{t}}\left(\sqrt{1-\beta_{t-1}}y_{t-2}+\gamma_{t-1}x+\sqrt{\beta_{t-1}}\epsilon_{t-1}\right)+\gamma_{t}x+\sqrt{\beta_{t}}\epsilon_{t}$ (23)
$=\sqrt{(1-\beta_{t})(1-\beta_{t-1})}y_{t-2}+\left(\sqrt{1-\beta_{t}}\gamma_{t-1}+\gamma_{t}\right)x+\text{noise terms}$
Continuing this expansion, we obtain:
$y_{t}=\left(\prod_{s=1}^{t}\sqrt{1-\beta_{s}}\right)y_{0}+\left(\sum_{s=1}^{t}\gamma_{s}\prod_{k=s+1}^{t}\sqrt{1-\beta_{k}}\right)x+\text{noise terms}$ (24)
Define $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$, then,
$y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\left(\sum_{s=1}^{t}\gamma_{s}\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}}\right)x+\text{noise terms}$ (25)
We define the cumulative bridge coefficient $\bar{\gamma}_{t}$ as
$\bar{\gamma}_{t}=\sum_{s=1}^{t}\gamma_{s}\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}}.$ (26)
The noise terms arise from the $\sqrt{\beta_{s}}\epsilon_{s}$ contributions at each step
$\text{noise terms}=$ $\sqrt{\beta_{t}}\epsilon_{t}+\sqrt{1-\beta_{t}}\sqrt{\beta_{t-1}}\epsilon_{t-1}+\sqrt{(1-\beta_{t})(1-\beta_{t-1})}\sqrt{\beta_{t-2}}\epsilon_{t-2}$ (27)
$+\cdots+\sqrt{\bar{\alpha}_{t}/\bar{\alpha}_{1}}\sqrt{\beta_{1}}\epsilon_{1}.$
This can be written compactly as:
$\text{noise terms}=\sum_{s=1}^{t}\left(\sqrt{\beta_{s}}\prod_{k=s+1}^{t}\sqrt{1-\beta_{k}}\right)\epsilon_{s}.$ (28)
Since the $\epsilon_{s}$ are independent, the total variance is the sum of the variances of each term:
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\beta_{s}\prod_{k=s+1}^{t}(1-\beta_{k}).$ (29)
we rewrite the variance:
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\beta_{s}\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}.$ (30)
Using $\bar{\alpha}_{s}=\prod_{k=1}^{s}(1-\beta_{k})$, we can express $\beta_{s}$ as $\beta_{s}=1-(1-\beta_{s})=1-\frac{\bar{\alpha}_{s}}{\bar{\alpha}_{s-1}}$. Substituting this in:
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\left(1-\frac{\bar{\alpha}_{s}}{\bar{\alpha}_{s-1}}\right)\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}.$ (31)
This simplifies to:
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s-1}}\right).$ (32)
This is a telescoping series:
$\text{Var}(\text{noise terms})=\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{1}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{0}}\right)+\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{2}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{1}}\right)+\cdots+\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}\right).$ (33)
Most terms cancel out, leaving:
$\text{Var}(\text{noise terms})=\bar{\alpha}_{t}\left(\frac{1}{\bar{\alpha}_{t}}-\frac{1}{\bar{\alpha}_{0}}\right).$ (34)
Assuming $\bar{\alpha}_{0}=1$ (since no steps have been applied at $t=0$):
$\text{Var}(\text{noise terms})=1-\bar{\alpha}_{t}.$ (35)
The variance of the accumulated noise is therefore $(1-\bar{\alpha}_{t})I$. Thus, Eq. (8) follows.
B.2 Proof of Proposition 4.2
We write $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. The bridge forward transition is
$q(y_{t}\mid y_{t-1},x)=\mathcal{N}\!\left(y_{t};\sqrt{\alpha_{t}}y_{t-1}+\gamma_{t}x,\beta_{t}I\right),$ (36)
and Proposition 4.1 gives
$q(y_{t-1}\mid y_{0},x)=\mathcal{N}\!\left(\sqrt{\bar{\alpha}_{t-1}}y_{0}+\bar{\gamma}_{t-1}x,(1-\bar{\alpha}_{t-1})I\right).$ (37)
The accumulated bridge coefficient satisfies the one-step recursion
$\bar{\gamma}_{t}=\gamma_{t}+\sqrt{\alpha_{t}}\bar{\gamma}_{t-1},$ (38)
which follows directly from $\bar{\gamma}_{t}=\sum_{s=1}^{t}\gamma_{s}\sqrt{\bar{\alpha}_{t}/\bar{\alpha}_{s}}$.
Gaussian posterior.
The exact bridge posterior is obtained from the product
$q(y_{t-1}\mid y_{t},y_{0},x)\propto q(y_{t}\mid y_{t-1},x)\,q(y_{t-1}\mid y_{0},x).$ (39)
Since both terms are Gaussian with isotropic covariance, the posterior is $\mathcal{N}(\tilde{\mu}_{t},\tilde{\beta}_{t}I)$, where
$\tilde{\beta}_{t}=\left(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}\right)^{-1}=\frac{\beta_{t}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}},$ (40)
and
$\tilde{\mu}_{t}=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})y_{t}\beta_{t}\sqrt{\bar{\alpha}_{t-1}}y_{0}\left(\beta_{t}\bar{\gamma}_{t-1}-\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t}\right)x}{1-\bar{\alpha}_{t}}.$ (41)
This is the standard Gaussian conditioning formula applied to the shifted linear observation $y_{t}=\sqrt{\alpha_{t}}y_{t-1}+\gamma_{t}x+\sqrt{\beta_{t}}\epsilon_{t}$.
Noise-prediction form.
Using the closed-form forward marginal,
$y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\qquad\epsilon\sim\mathcal{N}(0,I),$ (42)
we rewrite
$y_{0}=\frac{y_{t}-\bar{\gamma}_{t}x-\sqrt{1-\bar{\alpha}_{t}}\epsilon}{\sqrt{\bar{\alpha}_{t}}}.$ (43)
Substituting Eq. (43) into Eq. (41), the coefficients of $y_{t}$, $\epsilon$, and $x$ can be collected separately. For the $y_{t}$ coefficient,
$\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})+\beta_{t}\sqrt{\bar{\alpha}_{t-1}/\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}$ (44)
$\quad=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})+\beta_{t}/\sqrt{\alpha_{t}}}{1-\bar{\alpha}_{t}}=\frac{1}{\sqrt{\alpha_{t}}},$
where the last equality uses $\bar{\alpha}_{t}=\alpha_{t}\bar{\alpha}_{t-1}$. The noise coefficient becomes
$\frac{\beta_{t}\sqrt{\bar{\alpha}_{t-1}}\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}(1-\bar{\alpha}_{t})}=-\frac{\beta_{t}}{\sqrt{\alpha_{t}}\sqrt{1-\bar{\alpha}_{t}}}.$ (45)
Thus the only remaining non-standard term is the coefficient of $x$:
$C_{t}=\frac{-\beta_{t}\sqrt{\bar{\alpha}_{t-1}/\bar{\alpha}_{t}}\,\bar{\gamma}_{t}+\beta_{t}\bar{\gamma}_{t-1}-\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t}}{1-\bar{\alpha}_{t}}.$ (46)
Using $\sqrt{\bar{\alpha}_{t-1}/\bar{\alpha}_{t}}=1/\sqrt{\alpha_{t}}$ and the recursion in Eq. (38),
$-\frac{\beta_{t}}{\sqrt{\alpha_{t}}}\bar{\gamma}_{t}+\beta_{t}\bar{\gamma}_{t-1}$ (47)
$\quad=-\frac{\beta_{t}}{\sqrt{\alpha_{t}}}\left(\gamma_{t}+\sqrt{\alpha_{t}}\bar{\gamma}_{t-1}\right)+\beta_{t}\bar{\gamma}_{t-1}=-\frac{\beta_{t}}{\sqrt{\alpha_{t}}}\gamma_{t}.$
Therefore,
$C_{t}$ $=\frac{-\frac{\beta_{t}}{\sqrt{\alpha_{t}}}\gamma_{t}-\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t}}{1-\bar{\alpha}_{t}}$ (48)
$=-\frac{\gamma_{t}}{\sqrt{\alpha_{t}}}\frac{\beta_{t}+\alpha_{t}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}=-\frac{\gamma_{t}}{\sqrt{\alpha_{t}}},$
because $\beta_{t}+\alpha_{t}(1-\bar{\alpha}_{t-1})=1-\alpha_{t}\bar{\alpha}_{t-1}=1-\bar{\alpha}_{t}$.
Combining the three terms gives
$\tilde{\mu}_{t}=\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon\right)-\frac{\gamma_{t}}{\sqrt{\alpha_{t}}}x.$ (49)
Replacing the true noise $\epsilon$ by the learned predictor $\epsilon_{\theta}(y_{t},x,t)$ yields the parameterized reverse mean in the main text, with
$C_{t}(x)=-\frac{\gamma_{t}}{\sqrt{\alpha_{t}}}x=-\frac{\gamma_{t}}{\sqrt{1-\beta_{t}}}x.$ (50)
This simplification does not ignore the accumulated term $\bar{\gamma}_{t}x$; it uses the exact recursion in Eq. (38), so the accumulated bridge contribution cancels with the prior mean contribution in the Gaussian posterior. The result holds for any scalar bridge schedule $\{\gamma_{t}\}_{t=1}^{T}$ used in the forward transition; the SNR schedule is one particular choice.
B.3 Proof of Proposition 4.3
We derive the NBP loss in Proposition 4.3 directly from the Evidence Lower Bound (ELBO).
The log-likelihood of the data $y_{0}$ is lower-bounded by:
$\log p_{\theta}(y_{0}|x)\geq\mathbb{E}_{q(y_{1:T}|y_{0},x)}\left[\log\frac{p_{\theta}(y_{0:T}|x)}{q(y_{1:T}|y_{0},x)}\right]=\text{ELBO},$ (51)
where:
$p_{\theta}(y_{0:T}|x)$ is the reverse (generative) process.
$q(y_{1:T}|y_{0},x)$ is the forward (noising) process.
1. The ELBO decomposes into:
ELBO $=\mathbb{E}_{q(y_{1:T}|y_{0},x)}\left[\log p(y_{T}|x)+\sum_{t=2}^{T}\log\frac{p_{\theta}(y_{t-1}|y_{t},x)}{q(y_{t}|y_{t-1},x)}+\log\frac{p_{\theta}(y_{0}|y_{1},x)}{q(y_{1}|y_{0},x)}\right]$ (52)
$=\mathbb{E}_{q(y_{1:T}|y_{0},x)}\left[\log\frac{p(y_{T}|x)}{q(y_{T}|y_{0},x)}+\sum_{t=2}^{T}\log\frac{p_{\theta}(y_{t-1}|y_{t},x)}{q(y_{t-1}|y_{t},y_{0},x)}+\log p_{\theta}(y_{0}|y_{1},x)\right]$
The key term is the sum of KL divergences between $p_{\theta}(y_{t-1}|y_{t},x)$ and $q(y_{t-1}|y_{t},y_{0},x)$:
$\mathbb{E}_{q(y_{1:T}|y_{0},x)}\left[\sum_{t=2}^{T}D_{\text{KL}}(q(y_{t-1}|y_{t},y_{0},x)\|p_{\theta}(y_{t-1}|y_{t},x))\right]$ (53)
The KL divergence term between $q(y_{t-1}|y_{t},y_{0},x)$ and $p_{\theta}(y_{t-1}|y_{t},x)$ is:
$D_{\text{KL}}(q\|p_{\theta})\propto\|\mu_{\theta}(y_{t},t,x)-\tilde{\mu}(y_{t},y_{0},x)\|^{2}.$ (54)
where $\mu_{\theta}(y_{t},t,x)-\tilde{\mu}(y_{t},y_{0},x)$ are the means of $q(y_{t-1}|y_{t},y_{0},x)$ and $p_{\theta}(y_{t-1}|y_{t},x)$, respectively.
2. Sample from $q(y_{t}|y_{0},x)$: Using the forward marginal in Proposition 4.1, we can write:
$y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{I})$ (55)
This allows sampling $y_{t}$ directly from $y_{0}$, $x$ and $\epsilon$.
3. Rewrite $q(y_{t-1}|y_{t},y_{0},x)$ using Eq. (49):
$\tilde{\mu}_{t}(y_{t},y_{0},x)=\frac{1}{\sqrt{1-\beta_{t}}}\left(y_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon\right)+C(t)x.$ (56)
4. Reparameterize $\mu_{\theta}(y_{t},t,x)$: Assume $p_{\theta}(y_{t-1}|y_{t},x)$ predicts $\tilde{\mu}_{t}$:
$\mu_{\theta}(y_{t},t,x)=\frac{1}{\sqrt{1-\beta_{t}}}\left(y_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(y_{t},t,x)\right)+C(t)x$ (57)
Here, $\epsilon_{\theta}(y_{t},t,x)$ is a neural network predicting the noise $\epsilon$.
5. Final Noise-Prediction Loss: The KL terms simplify to a weighted $L_{2}$ loss on the noise:
$\mathbb{E}_{t,\epsilon}\left[\frac{\beta_{t}^{2}}{2\sigma_{t}^{2}\alpha_{t}(1-\bar{\alpha}_{t})}\|\epsilon-\epsilon_{\theta}(y_{t},t,x)\|^{2}\right]$ (58)
Dropping the weighting (as in DDPM) gives the simplified loss:
$\mathcal{L}=\mathbb{E}_{t,y_{0},x,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(y_{t},t,x)\|^{2}\right],$ (59)
where:
$t\sim\text{Uniform}(1,T)$,
$y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon$.
This proves Eq. (14), and the denoising network $\epsilon_{\theta}(y_{t},t,x)$ is self-consistent with respect to the condition on $x$.
B.4 Proof of Theorem 4.4
For fixed $y_{0}$, the two NBP marginals have the same covariance $(1-\bar{\alpha}_{t})I$ and means
$m_{t}(x)=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x,\qquad m_{t}(x^{\prime})=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x^{\prime}.$
For two Gaussians with identical covariance, the squared 2-Wasserstein distance equals the squared Euclidean distance between their means. Therefore,
$W_{2}^{2}=\|m_{t}(x)-m_{t}(x^{\prime})\|_{2}^{2}=\bar{\gamma}_{t}^{2}\|x-x^{\prime}\|_{2}^{2},$
which proves the claim. For NDP, the forward marginal does not contain $x$, so the two marginals are identical and the distance is zero.
B.5 Proof of Theorem 4.5
Conditioning on $y_{0}$, the NBP forward marginal can be written as a Gaussian channel
$y_{t}-\sqrt{\bar{\alpha}_{t}}y_{0}=\bar{\gamma}_{t}x+\eta_{t},\qquad\eta_{t}\sim\mathcal{N}(0,(1-\bar{\alpha}_{t})I),$
where $\eta_{t}$ is independent of $x$. If $x\sim\mathcal{N}(0,\Sigma_{x})$, the mutual information of this linear Gaussian channel is
$I(x;y_{t}\mid y_{0})=\frac{1}{2}\log\det\left(I+(1-\bar{\alpha}_{t})^{-1}\bar{\gamma}_{t}^{2}\Sigma_{x}\right).$
For NDP, $y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\eta_{t}$ conditional on $y_{0}$, so $y_{t}$ is independent of $x$ and the mutual information is zero.
B.6 Proof of Proposition 4.6
The NBP noised state is
$y_{t}(x)=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon.$
Applying the chain rule to $\ell_{t}(\theta,x)=\|\epsilon_{\theta}(y_{t}(x),x,t)-\epsilon\|_{2}^{2}$ gives
$\frac{d\ell_{t}}{dx}=\frac{\partial\ell_{t}}{\partial x}+\frac{\partial\ell_{t}}{\partial y_{t}}\frac{\partial y_{t}}{\partial x}=\frac{\partial\ell_{t}}{\partial x}+\bar{\gamma}_{t}\frac{\partial\ell_{t}}{\partial y_{t}}.$
The first term is the ordinary conditioning path through the denoiser input $x$, while the second term is induced by the input-anchored forward state. In NDP, the forward process does not contain $x$, so $\partial y_{t}/\partial x=0$ and the second term disappears.
Appendix C Background: Neural Processes
Neural Processes (NPs) [14, 13, 23, 28] combine the expressiveness of neural networks with the probabilistic reasoning of Gaussian Processes (GPs) [37]. While GPs offer principled uncertainty quantification, they suffer from poor scalability [42, 45] and limited kernel flexibility [46, 27]. In contrast, Neural Networks (NNs) [39, 33] are highly flexible and scalable but lack inherent mechanisms for uncertainty modeling [1, 35, 15]. NPs address these limitations by modeling distributions over functions using a neural network-based framework. They approximate a stochastic process $F:X\to Y$ through finite-dimensional marginals, parameterized by a latent variable $z$ to capture global uncertainty. Given context observations $(x_{\mathcal{C}},y_{\mathcal{C}})$ and target inputs $x_{\mathcal{T}}$, NPs generate predictive distributions over $y_{\mathcal{T}}$ via a conditional latent model.
$p(y_{\mathcal{T}},z|x_{\mathcal{T}},x_{\mathcal{C}},y_{\mathcal{C}})=p(z|x_{\mathcal{C}},y_{\mathcal{C}})\prod_{i=1}^{|\mathcal{T}|}p(y_{\mathcal{T},i}|x_{\mathcal{T},i},z)$ (60)
To ensure computational efficiency and order-invariance, NPs introduce three main components: 1. Encoder $h$: Maps each context input-output pair $(x_{i},y_{i})$ to a representation space, producing representations $r_{i}=h(x_{i},y_{i})$.
[…]
3. Decoder $g$: Predicts the target outputs $y_{\mathcal{T}}=g(x_{\mathcal{T}},z)$, conditioned on the latent variable $z$ and the target inputs $x_{\mathcal{T}}$.
Here, $z$ encodes the uncertainty about the global structure of the underlying function. Training NPs uses amortized variational inference, optimizing an evidence lower bound (ELBO) on the conditional log-likelihood:
$\log p(y_{\mathcal{T}}|x_{\mathcal{C}},y_{\mathcal{C}},x_{\mathcal{T}})\geq\mathbb{E}_{q(z|x_{\mathcal{C}},y_{\mathcal{C}},x_{\mathcal{T}},y_{\mathcal{T}})}\big[\sum_{i\in\mathcal{T}}\log p(y_{\mathcal{T},i}|z,x_{\mathcal{T},i})+\log\frac{p(z|x_{\mathcal{C}},y_{\mathcal{C}})}{q(z|x_{\mathcal{C}},y_{\mathcal{C}},x_{\mathcal{T}},y_{\mathcal{T}}))}\big]$ (61)
where $q(z|x_{\mathcal{C}},y_{\mathcal{C}},x_{\mathcal{T}},y_{\mathcal{T}})$ is the variational posterior distribution parameterized by a neural network, and $p(z|x_{\mathcal{C}},y_{\mathcal{C}})$ is the conditional prior.
Appendix D Input-output Dimensional Alignment
Real-world datasets often have mismatched input and output dimensions. Figure 3 summarizes how NBP handles this case. To apply the bridge in the output space, we introduce an anchor map
$a_{\psi}(x)=P_{\psi}(x),\qquad P_{\psi}:\mathbb{R}^{D_{x}}\rightarrow\mathbb{R}^{D_{y}},$ (62)
which maps each input location to an output-space anchor. The bridge term is then formed with $a_{\psi}(x)$ rather than the raw input $x$:
$q(y_{t}\mid y_{t-1},x)=\mathcal{N}\left(y_{t};\sqrt{1-\beta_{t}}y_{t-1}+\gamma_{t}a_{\psi}(x),\beta_{t}I\right).$ (63)
This design separates two roles of the input. The raw input $x$ is still passed to the denoising network as conditioning information, while $a_{\psi}(x)$ is used only to define an output-space bridge anchor. In image regression, for instance, $x$ is a two-dimensional normalized pixel coordinate and $y$ is a three-dimensional RGB value. We therefore parameterize $P_{\psi}$ as a lightweight coordinate-to-output anchor network: the coordinates are first positionally encoded, then passed through an MLP anchor head that outputs a vector in the RGB space. The anchor head is trained jointly with the denoising model using the main denoising objective and an auxiliary anchor reconstruction loss $\|a_{\psi}(x)-y_{0}\|^{2}$. This auxiliary term is used only to make the bridge anchor live in the correct output space; it does not change the denoising architecture, the number of reverse sampling steps, or the number of training updates. When $D_{x}=D_{y}$, as in the 1D GP experiments, $P_{\psi}$ can reduce to the identity map.
[…]
Figure 3: Learned input-output alignment for NBP when $D_{x}\neq D_{y}$. The raw input remains denoiser conditioning, while $P_{\psi}$ maps it into the output space before it enters the bridge trajectory. In image regression, this maps 2D pixel coordinates to RGB-valued anchors.
Appendix E NDP Review
Formally, given a function $f:\mathbb{R}^{D}\to\mathbb{R}$, an NDP learns a generative distribution over observed data pairs $(x,y)$, where inputs $x\in\mathbb{R}^{N\times D}$ and outputs $y=f(x)\in\mathbb{R}^{N}$. Unlike standard NPs, NDPs [10] do not explicitly require a partitioning into context and target sets during training; all points are jointly modeled. In supervised learning setting, the NDP modeling framework consists of two stochastic processes:
Forward Diffusion Process.
Starting from observed clean data $y_{0}$, the forward diffusion process gradually injects Gaussian noise into the outputs over $T$ timesteps according to a predefined variance schedule $\{\beta_{t}\}$:
$q(y_{1:T}\mid y_{0})=\prod_{t=1}^{T}q(y_{t}\mid y_{t-1}),\quad q(y_{t}\mid y_{t-1})=\mathcal{N}(y_{t};\sqrt{1-\beta_{t}}\,y_{t-1},\beta_{t}I).$ (64)
After $T$ diffusion steps, the distribution of the outputs converges towards standard Gaussian noise, i.e., $y_{T}\sim\mathcal{N}(0,I)$.
Reverse Process.
Neural Diffusion Processes (NDPs) learn a conditional reverse process that denoises observations from Gaussian noise $y_{T}$ to outputs $y_{0}$, guided by an input $x$:
$p_{\theta}(y_{0:T}\mid x)=p(y_{T})\prod_{t=1}^{T}p_{\theta}(y_{t-1}\mid y_{t},x),$ (65)
with Gaussian transitions parameterized by a noise prediction model $\epsilon_{\theta}$:
$p_{\theta}(y_{t-1}\mid y_{t},x)=\mathcal{N}\left(y_{t-1};\mu_{\theta}(y_{t},t,x),\tilde{\beta}_{t}I\right).$ (66)
where $\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$, $\mu_{\theta}(y_{t},t,x)$ is reparameterized as
$\mu_{\theta}(y_{t},t,x)=\frac{1}{\sqrt{1-\beta_{t}}}\left(y_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(y_{t},t,x)\right),\quad\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ (67)
Training Objective.
NDPs employ a denoising score matching objective  [22, 43, 21], training the noise model $\epsilon_{\theta}$ by minimizing the discrepancy between predicted noise and actual noise $\epsilon\sim\mathcal{N}(0,I)$:
$\mathcal{L}_{\theta}=\mathbb{E}_{t,x,y_{0},\epsilon}\left[\left\|\epsilon-\epsilon_{\theta}(y_{t},t,x)\right\|_{2}^{2}\right],\quad\text{with}\quad y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon.$ (68)
Conditional Sampling Procedure.
At test time, NDPs draw samples from the conditional distribution $p(y_{\mathcal{T},0}\mid x_{\mathcal{T}},D)$, where $D=(x_{\mathcal{C}}\in\mathbb{R}^{M\times D},y_{\mathcal{C},0}\in\mathbb{R}^{M})$ is the context observations.
The conditional sampling proceeds as follows. First, sample the initial target noise: $y_{\mathcal{T},T}\sim\mathcal{N}(0,I).$ For each diffusion timestep $t=T,\ldots,1$, proceed with:
  • Sample the noisy version of the context points using the forward diffusion process:
    $y_{\mathcal{C},t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}y_{\mathcal{C},0},(1-\bar{\alpha}_{t})I\right),$ (69)
  • Form the combined dataset at time $t$ by collecting the union of noisy targets and noisy contexts:
    $y_{t}=\{y_{\mathcal{T},t},y_{\mathcal{C},t}\},\quad x=\{x_{\mathcal{T}},x_{\mathcal{C}}\}.$ (70)
  • Perform the reverse denoising step by sampling from the learned backward kernel:
    $y_{t-1}\sim\mathcal{N}\left(\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(y_{t},t,x)\right),\sigma_{t}^{2}I\right),\quad\text{where}\quad\alpha_{t}=1-\beta_{t}.$ (71)
Method Endpoint Match Path Consistency
NDP (Baseline) Implicit Weak
NBP (Ours)
Table 6: Comparison of generation properties.
Appendix F Noise Model Architecture
To ensure that our model remains consistent with the structural properties of stochastic processes and to guarantee fair experimental comparisons, we adopt the same noise model architecture as NDPs, namely the Bi-Dimensional Attention Block [10].
Figure 4: Noise model architecture employed at each step of Neural Bridge Processes. The greyed-out box highlights the bi-dimensional attention block inherited from NDP.
As described in Figure 4, this architecture is designed to encode two key symmetries:
  • Exchangeability over data points: the model should be equivariant to permutations of the dataset ordering. That is, shuffling the order of inputs in the context or target set should not affect the output distribution.
  • Invariance over input dimensions: the prediction should be unaffected by reordering of input features (e.g., swapping the order of columns in a tabular dataset).
To accommodate both properties, the Bi-Dimensional Attention Block operates over a tensor $s_{t}\in\mathbb{R}^{N\times D\times H}$ representing the latent representation of paired inputs $(x,y_{t})$ and timestep $t$. Each block consists of two multi-head self-attention (MHSA) mechanisms:
  • $\text{MHSA}_{N}$: acts across the dataset axis $N$, propagating information across data points;
  • $\text{MHSA}_{D}$: acts across the input dimension axis $D$, capturing interactions between features.
The output of each block at layer $\ell$ is computed as:
$A_{t}^{(\ell)}(s_{t}^{(\ell-1)})=A_{t}^{(\ell-1)}+\sigma\left(\text{MHSA}_{N}(s_{t}^{(\ell-1)})+\text{MHSA}_{D}(s_{t}^{(\ell-1)})\right),$
where $\sigma$ denotes a ReLU activation, and $A_{t}^{(0)}=0$, $s_{t}^{(0)}=s_{t}$ is the output of the preprocessing stage.
Each Bi-Dimensional Attention Block maintains equivariance under permutations of data and feature dimensions:
Proposition F.1 (Equivariance [Prop. 4.1][10]).
Let $\pi_{N}$ and $\pi_{D}$ be permutations over dataset and feature axes respectively. Then,
$\pi_{D}\circ\pi_{N}\circ A_{t}(s)=A_{t}(\pi_{D}\circ\pi_{N}\circ s),\quad\forall s\in\mathbb{R}^{N\times D\times H}.$
The final noise model $\epsilon_{\theta}$ is obtained by summing outputs across all Bi-Attention layers, followed by an aggregation over the input dimension axis to remove dependence on feature order. This leads to:
Proposition F.2 ( Equivariance [Prop. 4.2][10]).
Let $\pi_{N},\pi_{D}$ be permutations as above. Then,
$\pi_{N}\circ\epsilon_{\theta}(x_{t},y_{t},t)=\epsilon_{\theta}(\pi_{N}\circ\pi_{D}\circ x_{t},\pi_{N}\circ y_{t},t).$
By encoding these structural properties, NDP and NBP generate, at each step, a set of exchangeable random variables— a structural requirement commonly used in NDP models to construct consistent finite-dimensional conditional distributions. However, as with NDP, our method does not attempt to satisfy the strict marginal consistency conditions required by the Kolmogorov Extension Theorem, even though the structure allows the model to empirically approximate consistency during meta-learning.
Appendix G More Details for Experiments
G.1 Baseline Implementation and Evaluation Metrics
To provide a comprehensive comparison, we implement NPs [14] , ANPs [23], and ConvNPs [16] using the official NP-Family repository [9], with all hyperparameters set to the recommended default values. For NDP [10], we directly utilize the official repositories. Meanwhile, to ensure a fair comparison, our NBP model adopts the same Bi-Dimensional Attention Block architecture and hyperparameters as the baseline NDP. The implementation details are provided in the supplementary materials. All models are retrained on the experimental datasets and evaluated with the same seed set to ensure consistent metric evaluation and visualization. We evaluate model performance using two primary metrics: Mean Squared Error (MSE) and Negative Log-Likelihood (NLL). Experiments are conducted on NVIDIA RTX A6000 GPUs with 48GB memory; each reported run uses a single GPU unless otherwise stated.
Figure 5: the signals recorded from these seven channels for a single trial of one subject.
G.2 Details of EEG Dataset Regression Task
The EEG dataset used in this experiment consists of recordings from 122 subjects, including both alcoholic and control groups. Each subject underwent either single or double stimulus conditions, during which neural responses were recorded using 64 scalp electrodes. Each trial lasted for 1 second with a sampling rate of 256 Hz, and up to 120 trials were recorded per subject.
For our study, we focus on signals from 7 frontal electrodes: FZ, F1, F2, F3, F4, F5, and F6. This selection yields a total of 7,632 multivariate time series, each comprising 256 time steps across 7 channels. These signals exhibit strong temporal dynamics and inter-channel correlations, making the dataset well-suited for evaluating the generalization and modeling capabilities of multi-output meta-learning models. The data is publicly available from the UCI Machine Learning Repository, with collection details described in [51]. Figure 5 illustrates the signals recorded from these seven channels for a single trial of one subject.
Subjects were split into training, validation, and test sets on a per-individual basis. The validation and test sets each contain 10 subjects, with the remainder assigned to the training set. All trials from each subject form a single meta-task, enabling task-level generalization evaluation.
Within each trial, we randomly select 3 out of the 7 channels and mask partial segments of these channels to simulate missing data. This setup supports three complementary prediction regimes. Interpolation evaluates whether the model can recover locally missing values within the observed time range. Reconstruction asks the model to infer masked regions of a target channel using the remaining channels as context, thereby testing cross-channel dependency modeling. Forecasting evaluates extrapolation to future signal trajectories based on currently observed measurements. Each input is represented as an index vector $\mathbf{x}_{e}=(i_{t},i_{c})$, where $i_{t}$ denotes the time step and $i_{c}$ the channel index. The corresponding output is the voltage signal $\mathbf{y}_{e}$.
All models were trained for 1,000 iterations on the training set. Evaluation metrics include Mean Squared Error (MSE) and Negative Log-Likelihood (NLL). The Neural Bridge Process (NBP) employs a 5-layer Bi-Dimensional Attention Block with hidden dimension 64 and 8 attention heads. We train and evaluate each model over multiple random seeds and report the mean performance, using the same seed set for all compared methods.
For models incorporating diffusion-based generation (such as NDP and NBP), we adopt a cosine noise schedule with the following parameters: $\beta_{\text{start}}=0.0003$, $\beta_{\text{end}}=0.5$, and 500 diffusion timesteps. These settings are applied consistently across the forward and reverse processes in all diffusion-based models. Additional training hyperparameters are as follows:
  • For NBP and NDP base, the learning rate was set to $2\cdot 10^{-5}$;
  • Other models used default learning rates as recommended in prior literature;
  • All models operated on input sequences of 256 time steps.
The evaluation results, summarized in Table 2 of the main text, show that NBP obtains lower NLL and MSE than the compared baselines across the three EEG prediction tasks.
Table 3: Calibration summary on the EEG interpolation setting. ECE and sharpness are estimated from predictive samples; lower ECE indicates better calibration, lower sharpness indicates more concentrated predictive uncertainty, and 90% coverage is best when close to 0.90.
Model NLL ($\downarrow$) MSE($\times 10^{-2}$) ($\downarrow$) ECE ($\downarrow$) Sharpness $\bar{\sigma}$ ($\downarrow$) 90% Coverage
NP 1.66 0.52 0.182 0.91 0.71
ANP 0.47 0.25 0.134 0.67 0.79
ConvNP 0.44 0.40 0.128 0.63 0.80
NDP -2.46 0.18 0.089 0.38 0.83
SNP -3.19 0.16 0.061 0.29 0.88
GEOMNDP -2.48 0.18 0.087 0.37 0.83
NBP (ours) -3.55 0.15 0.047 0.24 0.91
G.3 Details of Image Regression Task
(a) Ground Truth
(b) Corrupted Context
(c) Model Prediction
Figure 6: Qualitative results on the CelebA $\times64$ image regression task.
Table 7: CelebA 32 $\times$ 32 Results for NDP Base and NBP (Ours) in $10^{-2}$ MSE Units
Context Ratio Retained Pixels NDP Base NBP (Ours)
MSE Mean MSE Std MSE Mean MSE Std
0.1 96 1.7011 0.7221 1.4206 0.6958
0.2 197 0.8831 0.5214 0.7694 0.4008
0.3 312 0.4482 0.2724 0.4055 0.2157
0.4 412 0.2809 0.1511 0.2416 0.1091
0.5 512 0.2496 0.1251 0.2066 0.1062
0.6 570 0.1769 0.0793 0.1406 0.0573
0.7 714 0.1450 0.0729 0.1185 0.0579
0.8 832 0.0997 0.0518 0.0821 0.0373
0.9 913 0.0969 0.0527 0.0793 0.0392
Table 8: CelebA 64 $\times$ 64 Results for NDP Base and NBP (Ours) in $10^{-2}$ MSE Units
Context Ratio Retained Pixels NDP Base NBP (Ours)
MSE Mean MSE Std MSE Mean MSE Std
0.1 432 2.3763 1.3563 2.0189 1.2632
0.2 872 1.0546 0.6718 0.8615 0.5174
0.3 1276 0.7530 0.4283 0.6078 0.3132
0.4 1628 0.5115 0.2626 0.4432 0.1826
0.5 2080 0.4973 0.3144 0.4078 0.2548
0.6 2392 0.4160 0.2496 0.3587 0.1801
0.7 2788 0.3405 0.1985 0.3097 0.1694
0.8 3100 0.3215 0.1875 0.2838 0.1461
0.9 3492 0.2706 0.1579 0.2584 0.1177
We provide detailed information on the image regression task using Neural Bridge Processes (NBPs). The task involves predicting pixel intensities based on their spatial coordinates, which are normalized to the range [2,2][-2,2]. We use the CelebA dataset at resolutions of $32\times 32$ and $64\times 64$.
Our experimental protocol—including the denoising network architecture, training schedule, optimizer configuration, and seed set—closely follows the setup used for Neural Diffusion Processes (NDPs), ensuring a fair and consistent comparison. The core architecture of the NBP model consists of 7 layers, each with hidden dimension 64 and 8 attention heads. Sparse attention is not used in these experiments.
Training Configuration.
The model is trained using a batch size of 32. The optimizer follows the same learning-rate schedule as the NDP baseline:
  • Initial learning rate: $2.0\times 10^{-5}$
  • Peak learning rate: $1.0\times 10^{-3}$
  • End learning rate: $1.0\times 10^{-5}$
  • Warmup and cosine decay are applied following the NDP implementation.
  • EMA decay rate: 0.995
Diffusion Settings.
We employ a cosine beta schedule with the following parameters for the forward and reverse processes:
  • $\beta_{\text{start}}=0.0003$
  • $\beta_{\text{end}}=0.5$
  • Number of timesteps: 500
Evaluation Protocol.
Each prediction is averaged over 9 conditional samples during testing. The evaluation batch size is set to 9, with 128 samples drawn per image for final averaging. All pixel values are normalized to the $[0,1]$ range. Results are averaged over multiple random seeds using the same seed set across compared methods.
Loss Function.
We adopt the $\ell_{1}$ loss for training the denoising objective.
Results and Analysis.
Tables 7 and 8 report the quantitative performance under various levels of context sparsity. NBPs obtain lower MSE than NDPs across the reported settings. For example, at a context ratio of 0.2 on CelebA $32\times 32$, NBPs achieve an MSE of 0.76 compared to 0.88 by NDPs. This trend persists at higher resolutions: on CelebA $64\times 64$, NBPs achieve an MSE of 0.86 compared to NDPs’ 1.05 under the same sparse context condition.
These results are consistent with the proposed design, in which the forward diffusion kernel is explicitly conditioned on the input coordinates through the bridge term. The component ablation in Appendix H.2 further separates this effect from endpoint initialization, learned coordinate-anchor capacity, and denoiser capacity.
Appendix H Ablation Study
H.1 Bridge Coefficient Schedules
To validate the design of our bridge coefficient $\gamma_{t}$, we conduct an ablation study comparing three scheduling strategies on CelebA $32\times 32$ image regression. All experiments use identical architecture (7-layer Bi-Dimensional Attention Block, 64 hidden dimensions, 8 attention heads) and training configuration (batch size 32, cosine beta schedule with $\beta_{\text{start}}=0.0003$, $\beta_{\text{end}}=0.5$, 500 timesteps).
Bridge Coefficient Schedules.
We compare the following three strategies for computing $\gamma_{t}$:
  • Linear Schedule: $\gamma_{t}=\frac{t}{T}$, where $T=500$ is the total number of timesteps. This provides a uniform increase in the influence of $x$ over time.
  • Cosine Schedule: $\gamma_{t}=\frac{1-\cos(\pi t/T)}{2}$. This schedule starts slowly, accelerates in the middle phase, and saturates towards the end.
  • SNR-based Schedule (Ours): $\gamma_{t}=\frac{\text{SNR}_{T}}{\text{SNR}_{t}}$, where $\text{SNR}_{t}=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. This design explicitly couples the bridge strength to the signal-to-noise ratio, ensuring theoretical consistency with the diffusion dynamics.
Evaluation Protocol.
We evaluate reconstruction performance at nine context ratios (0.1 to 0.9) on the CelebA test set. For each method, we report Mean Squared Error (MSE) averaged over 128 test images, with each prediction obtained by averaging 9 conditional samples.
Results.
Table 9 presents the quantitative comparison. The SNR-based schedule consistently outperforms both linear and cosine schedules across all context ratios, with the largest improvements observed in low-data regimes (context ratio 0.1–0.3). For instance, at context ratio 0.1, our SNR-based method achieves MSE of $1.4206\times 10^{-2}$ compared to $1.6823\times 10^{-2}$ (linear) and $1.5491\times 10^{-2}$ (cosine), representing relative improvements of 15.6% and 8.3% respectively.
Figure 7 visualizes the behavior of the three schedules over the diffusion trajectory. The SNR-based schedule maintains lower bridge coefficients during early timesteps, where the clean signal remains relatively strong, and increases the bridge weight during later timesteps, where noise dominates. This behavior is consistent with the observed improvement of the SNR schedule in Table 9.
Table 9: Ablation study on CelebA $32\times 32$: Comparison of bridge coefficient schedules. MSE values are reported in $10^{-2}$ units. Best results are bolded.
Context Ratio Linear Cosine SNR (Ours)
0.1 1.6823 $\pm$ 0.7458 1.5491 $\pm$ 0.7185 1.4206 $\pm$ 0.6958
0.2 0.9247 $\pm$ 0.5421 0.8312 $\pm$ 0.4687 0.7694 $\pm$ 0.4008
0.3 0.4891 $\pm$ 0.2965 0.4428 $\pm$ 0.2441 0.4055 $\pm$ 0.2157
0.4 0.2975 $\pm$ 0.1673 0.2654 $\pm$ 0.1298 0.2416 $\pm$ 0.1091
0.5 0.2698 $\pm$ 0.1392 0.2331 $\pm$ 0.1184 0.2066 $\pm$ 0.1062
0.6 0.1924 $\pm$ 0.0891 0.1632 $\pm$ 0.0698 0.1406 $\pm$ 0.0573
0.7 0.1589 $\pm$ 0.0798 0.1364 $\pm$ 0.0672 0.1185 $\pm$ 0.0579
0.8 0.1098 $\pm$ 0.0587 0.0947 $\pm$ 0.0456 0.0821 $\pm$ 0.0373
0.9 0.1065 $\pm$ 0.0594 0.0912 $\pm$ 0.0462 0.0793 $\pm$ 0.0392
Figure 7: Visualization of the three bridge coefficient schedules over 500 diffusion timesteps. The SNR-based schedule (orange) adaptively increases the influence of input $x$ based on the signal-to-noise ratio, providing stronger guidance in later timesteps where noise dominates.
Figure 8: Reconstruction error comparison across different context ratios. In this ablation, the SNR-based schedule achieves the lowest MSE, with larger observed gains in sparse context regimes (context ratio $\leq$ 0.3).
Analysis.
The SNR-based schedule is motivated by matching the bridge strength to the diffusion signal-to-noise ratio. Unlike the linear and cosine schedules tested here, it adjusts the bridge coefficient $\gamma_{t}$ according to the diffusion process’s intrinsic dynamics. Specifically, by setting $\gamma_{t}=\text{SNR}_{T}/\text{SNR}_{t}$, the schedule has the following properties:
  1. 1.
    Early phase consistency: When $t\ll T$, $\text{SNR}_{t}\gg\text{SNR}_{T}$, leading to $\gamma_{t}\to 0$. This allows the process to approximate standard diffusion, preserving the data distribution.
  2. 2.
    Endpoint anchoring: As $t\to T$, $\text{SNR}_{t}\to\text{SNR}_{T}$, driving $\gamma_{t}\to 1$. This increases the input-dependent shift in the terminal region of the forward process.
This adaptive behavior provides a simple way to balance early-stage data preservation with later-stage input anchoring, which is consistent with the lower reconstruction errors observed under sparse context conditions.
Computational Cost.
Importantly, all three schedules incur identical computational cost during both training and inference, as the schedule computation is a simple preprocessing step that does not affect the denoising network architecture or the number of diffusion steps. Therefore, this ablation isolates the schedule choice from model capacity and sampling budget.
H.2 Component Ablation on CelebA NLL
To clarify where the improvements in Table 5 come from, we further ablate the main components of NBP on CelebA image regression. All variants use the same evaluation protocol, denoising backbone, optimizer, number of diffusion steps, and training budget unless explicitly stated. We report NLL under half-image and random-context observations. Lower values are better. The last column shows the absolute improvement relative to the NDP baseline.
Table 10: Component ablation for CelebA image regression NLL. Lower is better. The relative improvement is computed against the NDP baseline. The fixed-$P$ row removes the learned anchor parameters, while the strong-denoiser row controls for increased denoising capacity without using a bridge.
Method Half NLL Random NLL $\Delta$ vs. NDP
NDP (baseline) -3.92 -3.98
Initialization only -4.13 -4.22 -0.21 / -0.24
Bridge only -4.08 -4.16 -0.16 / -0.18
Full + fixed $P$ -4.71 -4.93 -0.79 / -0.95
Full + learned $P_{\psi}$ -4.83 -5.05 -0.91 / -1.07
SNP strong denoiser, no bridge -4.76 -4.77 -0.84 / -0.79
NBP (ours) -5.02 -5.24 -1.10 / -1.26
The ablation shows that neither endpoint initialization nor the bridge term alone explains the full improvement. Both components help, but the largest gains arise when the complete bridge construction is combined with output-space alignment. The fixed-$P$ variant keeps the bridge mechanism but removes the learned anchor parameters and auxiliary anchor reconstruction loss; its improvement over NDP indicates that the bridge itself contributes substantially. The learned-$P_{\psi}$ variant then isolates the additional benefit of adapting the anchor to the output space. Finally, the SNP strong-denoiser baseline improves denoising capacity without using an input-anchored forward path, but it remains worse than NBP. These controls suggest that the gain is not solely due to extra parameters, auxiliary supervision, or a stronger denoising network; the process-level bridge provides a complementary effect. This is not meant to claim perfect parameter-count equality for every variant, but to separate backbone capacity, anchor learning, and path-level bridging as much as possible under a shared training protocol.
H.3 Transferring the Bridge Idea to Flow Matching
The main NBP formulation instantiates the bridge idea in a DDPM-style diffusion process. To test whether the benefit is specific to diffusion or instead comes from the more general principle of input-anchored generative paths, we conduct an additional controlled ablation using Flow Matching Neural Processes (FlowNP) [17]. FlowNP learns a velocity field along a continuous probability path from a base sample $y_{0}$ to the observed target value $y_{1}$. In the standard implementation, this path is input-independent:
$y_{t}=(1-t)y_{0}+ty_{1},\qquad v_{t}=y_{1}-y_{0}.$ (72)
We construct a bridge-augmented variant, denoted FBP, by anchoring the intermediate states to the target input $x$:
$y_{t}=(1-t)y_{0}+ty_{1}+\gamma(t)x,\qquad v_{t}=y_{1}-y_{0}+\gamma^{\prime}(t)x.$ (73)
For this ablation, we use a smooth bump schedule $\gamma(t)=4\lambda t(1-t)$, so $\gamma(0)=\gamma(1)=0$. Therefore the source and target endpoints are unchanged, while only the intermediate generative trajectory is modified. This is important because it isolates the effect of path-level input anchoring without changing the target distribution or the likelihood evaluation protocol.
Experimental Setup.
We evaluate on the 1D Gaussian-process regression benchmark with an RBF kernel. FBP uses exactly the same transformer backbone, positional encoding, optimizer, number of training steps, and ODE likelihood estimator as FlowNP. Both FBP and FlowNP have 224,449 trainable parameters. Thus, the comparison isolates the bridge path design rather than model capacity. We report target log-likelihood (LL, higher is better) at training steps 5k, 10k, 15k, and 20k. We also include NDP as a diffusion-based reference.
Figure 9: Posterior samples on a 1D GP task using the same context set. FBP uses the FlowNP backbone but replaces the standard input-independent flow path with an input-anchored bridge path.
Table 11: Bridge-transfer ablation on 1D GP regression with RBF kernel. FBP applies the input-anchored bridge path to the FlowNP framework while keeping the same architecture and ODE likelihood evaluation. Target $L$ is reported; higher is better.
Method 5k 10k 15k 20k
NDP -1.8140 0.0186 0.5407 0.6331
FlowNP -0.4501 1.0328 1.3558 1.4215
FBP, $\lambda=0.1$ 1.1374 1.3959 1.5480 1.5857
FBP, $\lambda=0.2$ 1.2256 1.3832 1.5298 1.5937
FBP, $\lambda=0.3$ 1.1706 1.4289 1.5529 1.5878
FBP, $\lambda=0.5$ 1.1002 1.4002 1.5324 1.5805
Analysis.
The bridge-augmented flow path consistently improves over the FlowNP baseline across all tested bridge strengths. At 20k steps, the best FBP variant achieves target LL 1.5937, compared with 1.4215 for FlowNP and 0.6331 for NDP. The improvement is also robust to the bridge strength $\lambda$: all tested FBP variants reach 20k LL above 1.58. These results suggest that the benefit of NBP is not tied only to DDPM-style diffusion. Rather, input anchoring is a more general path-design principle: injecting input information into the generative trajectory itself can strengthen conditional function modeling even when the underlying generative mechanism is flow matching instead of diffusion.
H.4 Quantifying Path Supervision: Sensitivity Analysis
A key advantage of NBP over NDP lies in its input-anchored forward kernel (Eq. 8 and 10 in the main text), which explicitly injects input $x$ into the entire diffusion trajectory rather than relying solely on the denoiser to implicitly learn conditional guidance. To quantify this advantage, we conduct a sensitivity analysis measuring how strongly the denoiser $\epsilon_{\theta}(y_{t},x,t)$ depends on the input $x$ at different timesteps.
Measurement: Jacobian Sensitivity.
For a given sample, we compute the sensitivity of the denoiser output with respect to input $x$ at each timestep $t$:
$S(t)=\mathbb{E}\left[\left\|\frac{\partial\epsilon_{\theta}(y_{t},x,t)}{\partial x}\right\|_{F}\right]$ (74)
where $\|\cdot\|_{F}$ denotes the Frobenius norm. In PyTorch, we enable requires_grad=True for $\epsilon_{\theta}$, compute the norm of gradients with respect to $x$, and then normalize both the output gradient and input gradient to obtain relative sensitivity. This metric directly measures the effective coupling between input and output through the learned denoiser.
Hypothesis.
We expect the following behavior:
  • NDP: $S(t)$ should increase gradually from early to late timesteps, approaching 0 in early phases (because the forward process does not contain $x$, relying solely on implicit "soft conditioning" through the denoiser).
  • NBP: $S(t)$ should remain substantial across all timesteps, especially in early phases, reflecting the explicit injection of $x$ into the forward kernel via $\gamma_{t}x$.
Experimental Setup.
We evaluate on CelebA $32\times 32$ with context ratio 0.3. For each method (NDP and NBP), we:
  1. 1.
    Sample 100 test images and randomly mask 70% of pixels.
  2. 2.
    For each timestep $t\in\{1,50,100,150,\ldots,500\}$, forward diffuse the image to obtain $y_{t}$.
  3. 3.
    Compute $S(t)$ by averaging the Jacobian norm $\|\partial\epsilon_{\theta}/\partial x\|_{F}$ over all samples.
Results.
[…]
Additional Metrics: Gradient Statistics.
To further characterize training dynamics, we also measure:
  • Gradient magnitude: $G(t)=\mathbb{E}[\|\nabla_{\theta}\mathcal{L}(t)\|_{2}]$, the average gradient norm with respect to denoiser parameters at timestep $t$.
  • Gradient variance: $\text{Var}(t)=\text{Var}[\|\nabla_{\theta}\mathcal{L}(t)\|_{2}]$, measuring training stability.
Table 12 reports these statistics across three representative phases: early ($t\in[1,100]$), middle ($t\in[200,300]$), and late ($t\in[400,500]$). In this setting, NBP exhibits larger average gradient magnitudes and lower gradient variance, with the clearest difference in the early phase. These measurements are consistent with the direct-gradient analysis in Proposition 4.6.
Table 12: Gradient statistics during training on CelebA $32\times 32$ with context ratio 0.3. Values are diagnostic estimates over representative timestep ranges.
Phase Method Grad. Magnitude $\uparrow$ Grad. Variance $\downarrow$
Early ($t\leq 100$) NDP $2.43\times 10^{-3}$ $1.86\times 10^{-6}$
NBP $\mathbf{3.45\times 10^{-3}}$ $\mathbf{1.28\times 10^{-6}}$
Middle ($200\leq t\leq 300$) NDP $4.12\times 10^{-3}$ $2.31\times 10^{-6}$
NBP $\mathbf{4.89\times 10^{-3}}$ $\mathbf{1.94\times 10^{-6}}$
Late ($t\geq 400$) NDP $5.67\times 10^{-3}$ $3.45\times 10^{-6}$
NBP $\mathbf{6.21\times 10^{-3}}$ $\mathbf{2.98\times 10^{-6}}$
Interpretation.
The elevated sensitivity $S(t)$ in NBP’s early timesteps directly stems from the explicit $\gamma_{t}x$ term in the forward kernel (Eq. 8). During training, the loss $\mathcal{L}=\mathbb{E}[\|\epsilon-\epsilon_{\theta}(y_{t},x,t)\|^{2}]$ with $y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ (Eq. 10) propagates gradients not only through the denoiser $\epsilon_{\theta}$ but also through the $\bar{\gamma}_{t}x$ term embedded in $y_{t}$. This creates a dual pathway for supervision:
  1. 1.
    Direct path: $x\to y_{t}\to\epsilon_{\theta}$ (via forward kernel)
  2. 2.
    Indirect path: $x\to\epsilon_{\theta}$ (via denoiser conditioning)
In contrast, NDP relies entirely on the indirect path, placing the full burden on the denoiser-side conditioning mechanism. The sensitivity analysis is consistent with the presence of this additional direct path in NBP, especially when the signal-to-noise ratio is high (early $t$).
Implications.
These diagnostics suggest several possible explanations for the empirical improvements:
  • Optimization: Larger early-phase gradients may make input-output dependencies easier to learn.
  • Trajectory conditioning: Input-dependent training states reduce the reliance on the denoiser-side conditioning pathway alone.
  • Prediction quality: The bridge path provides an additional mechanism for aligning generated outputs with the observed input structure.
In summary, the sensitivity analysis provides diagnostic evidence that NBP changes where input information enters the diffusion model. We do not treat this metric as a direct performance measure; rather, it complements the quantitative results and ablations by illustrating the proposed path-level conditioning mechanism.
Appendix I Code Contribution
The full implementation of the Neural Bridge Processes (NBP) framework is provided in the supplementary materials to ensure reproducibility and to facilitate further evaluation by reviewers.
Appendix J Statement on the Use of Large Language Models
Large language models (LLMs) were used solely for polishing and editing the text of this manuscript.
Appendix K Limitations
NBP is most effective when the input provides a meaningful anchor for the output trajectory. If the input and output spaces are weakly aligned, or if the learned anchor map used to match their dimensions is poorly estimated, the bridge term can provide a misleading endpoint signal. This limitation is especially relevant when $D_{x}\neq D_{y}$, where the method relies on the output-space anchor $a_{\psi}(x)=P_{\psi}(x)$.
As with other diffusion-based stochastic process models, sampling cost scales with the number of denoising steps. Although the bridge coefficients and reverse correction introduce little computational overhead relative to NDP, very high-dimensional outputs may still require additional architectural or schedule design. Our current evaluation focuses on synthetic regression, EEG regression, CylinderFlow, image regression, and a bridge-transfer study on 1D GP regression with FlowNP. Broader domains, including spatiotemporal modeling, control, scientific data analysis, and more challenging high-dimensional output spaces, remain important directions for future work.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Guidelines:
    • The answer [N/A] means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A [No] or [N/A] answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Guidelines:
    • The answer [N/A] means that the paper has no limitation while the answer [No] means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate “Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Guidelines:
    • The answer [N/A] means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • If the paper includes experiments, a [No] answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Guidelines:
    • The answer [N/A] means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so [No] is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer) necessary to understand the results?
  23. Answer: [Yes]
  24. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • The authors should answer [Yes] if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
    • If error bars are reported in tables or plots, the authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Guidelines:
    • The answer [N/A] means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer [No] , they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [N/A]
  40. Guidelines:
    • The answer [N/A] means that there is no societal impact of the work performed.
    • If the authors answer [N/A] or [No] , they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate Deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pre-trained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Guidelines:
    • The answer [N/A] means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Guidelines:
    • The answer [N/A] means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Guidelines:
    • The answer [N/A] means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [Yes]
  56. Guidelines:
    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Guidelines:
    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigor, or originality of the research, declaration is not required.
  63. Answer: [Yes]
  64. Guidelines:
    • The answer [N/A] means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy in the NeurIPS handbook for what should or should not be described.